OpenStudy (anonymous):

Solve for x in the proportion 3x/3x-36 = -1/x

OpenStudy (anonymous):

Ok...u can cross multiply to solve this problem so... it will look like 3x(x) = -1(3x-36) then, it will be 3x^2 = -3x + 36 and now, its 3x^2 + 3x -36 = 0 Its basically a factoring problem from here...can you solve it from here?

OpenStudy (anonymous):

uhhh, possibly..

OpenStudy (anonymous):

ok...what would be your first step?

OpenStudy (anonymous):

take out a 3?

OpenStudy (anonymous):

sweet! do it!:)

OpenStudy (anonymous):

okay 3(x^2+x-12)

OpenStudy (anonymous):

right! 3(x^2 + x - 12) = 0 then, (x^2 + x - 12) = 0/3 = 0 now, can you factor (x^2 + x - 12) = 0

OpenStudy (anonymous):

um, well what would i do?

OpenStudy (anonymous):

ok...i'll just do it! you can use simple factoring to see that (x^2 + x - 12) factors into (x + 4) (x - 3) = 0 then, you have (x + 4) = 0 x = -4 and (x - 3) = 0 x = 3 These are your values for x!

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

my pleasure!

OpenStudy (yolo_king):

I GOT A HUNDRED PERCENT ON MY ASSIGNMENT

OpenStudy (yolo_king):

oops sry wrong post